OpenStudy (anonymous):

At Sami's Shoe Warehouse, it takes fraction 4 over 5 of a day to complete fraction 1 over 10 of an order of sneakers. At this rate, how long will it take to complete the entire order of sneakers? 6 days 8 days Mixed number 11 and 1 over 4 days Mixed number 12 and 1 over 2 days

OpenStudy (anonymous):

@Michele_Laino pls help

OpenStudy (michele_laino):

I think are necessary \[\frac{ 4 }{ 5 }*10=8 days\]

OpenStudy (anonymous):

so b, thank you!!!!!!!!!!!

OpenStudy (anonymous):

An ice sculpture is melting because of heat. Assuming its height changes by negative 2 over 25 m every hour, what will be the change in the height after 5 hours? negative 2 over 5 meter negative 2 over 125 meter 2 over 125 meter 2 over 5 meter

OpenStudy (michele_laino):

yes, b! thank you!!!!!

OpenStudy (anonymous):

youre welcome:D

OpenStudy (michele_laino):

thanks, you also are welcome!

OpenStudy (anonymous):

what about the second one??

OpenStudy (michele_laino):

I think the change is: \[\frac{ 2 }{ 25 }*5=\frac{ 2 }{ 5 }\] so the last option

OpenStudy (anonymous):

thx

OpenStudy (michele_laino):

thanks!

OpenStudy (anonymous):

Joan has red and pink hibiscus in groups of 9. She has x groups of red hibiscus and y groups of pink hibiscus. Which expression shows the total number of hibiscus that Joan has? 9x + y x + 9y 9x + 9y x over 9 + y over 9

OpenStudy (michele_laino):

I think 9x+9y, because in each group there are 9 hibiscus

OpenStudy (anonymous):

thx

OpenStudy (michele_laino):

thanks!

OpenStudy (anonymous):

Simplify (−8.5)(−5)( −2). 850 85 −85 −850

OpenStudy (anonymous):

-

OpenStudy (michele_laino):

we have: \[(-8.5)(5)(2)=(-8.5)[(5)(2)]=8-8.5)[10]=-85\]

OpenStudy (anonymous):

so c??

OpenStudy (michele_laino):

please I rewrite:

OpenStudy (anonymous):

Diana made a recipe that yields 6 and 1 over 2 cups. If each serving is 1 over 4 cup, which expression will help Diana determine the number of servings her recipe will yield? 6 and 1 over 2 ⋅ 1 over 4 6 and 1 over 2 ÷ 1 over 4 6 and 1 over 2 + 1 over 4 6 and 1 over 2 − 1 over 4

OpenStudy (michele_laino):

\[(-8.5)(-5)(-2)=(-8.5)[(-5)(-2)]=(-8.5)[10]=-85\] so c.

OpenStudy (anonymous):

\[1\times2\div3+4\]

OpenStudy (anonymous):

#

OpenStudy (anonymous):

4 is minus

OpenStudy (michele_laino):

I think the second option, namely: 6 and 1 over 2 ÷ 1 over 4

OpenStudy (anonymous):

thank you very much @Michele_Laino

OpenStudy (anonymous):

Remo needs to buy a countertop for a bathroom. He calculated the area as 10 square feet. The actual area is 9.5 square feet. What is Remo's percentage error? 0.5% 0.95% 5% 5.26%

OpenStudy (anonymous):

c??

OpenStudy (michele_laino):

\[1\times 2 \div3 +4=2 \div3+4=4.66\]

OpenStudy (anonymous):

so then it would be c?

OpenStudy (anonymous):

Simplify 5 − (−1). 6 −6 4 −4

OpenStudy (anonymous):

-

OpenStudy (michele_laino):

I think (10-9.5)/9.5=5.26%

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

so simplify 5 - (-1)

OpenStudy (anonymous):

6, -6, 4, -4

OpenStudy (michele_laino):

5-(-1)=5+1=6, so A

OpenStudy (anonymous):

thx

OpenStudy (michele_laino):

thanks!

OpenStudy (anonymous):

Highlighter pens cost $3.86 per set at an office supply store. Which equation can be used to determine the price, P, of n sets of these pens? n = 3.86P P equals 3.86 over n n equals 3.86 over P P = 3.86n

OpenStudy (michele_laino):

I think that your answer is: P=3.86*n, so the last option

OpenStudy (anonymous):

thx :P

OpenStudy (michele_laino):

thanks!:)